### v1

# TOTEM: TOkenized Time Series EMbeddings for General Time Series Analysis

Abstract
The field of general time series analysis has recently begun to explore unified modeling, where a common architectural backbone can be retrained on a specific task for a specific dataset. In this work, we approach unification from a complementary vantage point: unification across tasks and domains. To this end, we explore the impact of discrete, learnt, time series data representations that enable generalist, cross-domain training. Our method, TOTEM, or TOkenized Time Series EMbeddings, proposes a simple tokenizer architecture that embeds time series data from varying domains using a discrete vectorized representation learned in a self-supervised manner. TOTEM works across multiple tasks and domains with minimal to no tuning. We study the efficacy of TOTEM with an extensive evaluation on 17 real world time series datasets across 3 tasks. We evaluate both the specialist (i.e., training a model on each domain) and generalist (i.e., training a single model on many domains) settings, and show that TOTEM matches or outperforms previous best methods on several popular benchmarks. The code can be found at: https://github.com/SaberaTalukder/TOTEM.
representation learning, tokenization, time series
1 Introduction
Figure 1: TOTEM & Evaluation Regimes. An overview of TOTEM, the training schemas, and inference regimes. (a) The TOTEM VQVAE architecture consists of an 1D strided CNN encoder $\mathcal{E}$, quantizer, latent codebook, and 1D strided transpose CNN decoder $\mathcal{D}$. TOTEM’s VQVAE enables generalist training, i.e., on all datasets jointly, and specialist training, i.e., on one dataset at a time. (b) TOTEM’s discrete, self-supervised codebook can be leveraged for both in domain and zero shot testing.
Time series analysis encompasses a wide range of tasks, datasets, and applications in the real world. When considering training paradigms, time series analysis has historically been conducted via specialist-training, meaning that models are trained on a single time series domain (Zhou et al., 2023; Wu et al., 2022a; Nie et al., 2022; Zhang & Yan, 2022). Generalist-training, where models are simultaneously trained on multiple time series domains, contrasts the specialist paradigm. Both specialist and generalist models can be tested under various regimes. Within in-domain-testing, a model is tested on the same domain(s) it was trained on. In zero-shot-testing, a model is tested on different domains(s) than it was trained on. Some methods have begun to explore the idea of zero-shot forecasting where (1) a forecaster trains on one dataset then predicts on a separate dataset (Zhou et al., 2023), or (2) a forecaster trains on a subset of channels (which we call sensors) from one dataset then zero-shot forecasts on the remaining sensors in the same dataset (Liu et al., 2023). Both of these models would be considered specialists, as they were trained on only one (or a subset of one) dataset.
Further, time series analysis has typically been restricted by task, where methods study only forecasting (Wu et al., 2021; Woo et al., 2022), anomaly detection (Xu et al., 2021; He & Zhao, 2019), or imputation (Luo et al., 2018, 2019; Talukder et al., 2022), among others. Recently, the field has become increasingly unified with respect to model architecture, with methods (Zhou et al., 2023; Wu et al., 2022a) exploring language and vision backbones on various time series tasks. These backbones, like previous methods, utilize specialist training (e.g., training separate imputers on each dataset).
Similarly, the field has also become increasingly unified with respect to data representation, with growing emphasis on learning performant data representations. For instance, Franceschi et al. (2019) utilize an exponentially dilated causal convolutional encoder to discover in-domain embeddings, Tonekaboni et al. (2021) leverage temporal neighborhood coding, Yang & Hong (2022) utilize temporal-spectral fusion, and Yue et al. (2022) employs hierarchical contrasting across the time and batch dimensions.
At a technical level, our approach bears closest affinity to methods that use vector quantized variational autoencoders (VQVAEs) (Van Den Oord et al., 2017; Duan et al., 2023; Rasul et al., 2022b, a). As we discuss further in Section 3, our goal is to develop a streamlined framework for learning a tokenized data representation (using VQVAEs) in a way that permits easy applicability and holistic empirical evaluation on a broad range of time series modeling tasks and data domains (including zero-shot generalization to new test domains) with minimal to no tuning. [^2] As an aside, our approach to studying what is a performant general time series data representation shares a philosophical alignment with the development of large generalist models in natural language processing, which are also based on having a common tokenized representation (Gage, 1994; Radford et al., 2018).
Motivated by the trend of time series analysis unification, we explore the value of a VQVAE-based tokenizer for time series imputation, anomaly detection, and forecasting. Unlike previous methods, we utilize self-supervised, discrete tokens, and extensively explore their utility in varied training and testing regimes. Our contributions are as follows:
  1. 1.
    We present TOTEM, a simple tokenizer architecture for time series analysis that works across domains and tasks with minimal to no tuning.
  2. 2.
    Despite its simplicity, TOTEM matches or outperforms the state-of-the-art on several popular benchmark datasets and tasks.
  3. 3.
    With an extensive evaluation in the generalist setting (training a single model on multiple domains), we show that TOTEM outperforms the leading state-of-the-art model in both in-domain and zero-shot testing regimes.
2 Related Work
Time series modeling methods utilize many techniques, ranging from statistical methods (Winters, 1960; Holt, 1957; Anderson, 1976; Hyndman & Athanasopoulos, 2018; Taylor & Letham, 2018) to multilayer perceptrons (MLPs) (Zeng et al., 2023; Li et al., 2023; Das et al., 2023; Challu et al., 2023; Chen et al., 2023; Zhang et al., 2022; Oreshkin et al., 2019) to convolutional neural networks (CNNs) (Wu et al., 2022a; Liu et al., 2022a; He & Zhao, 2019; Franceschi et al., 2019; Bai et al., 2018) to recurrent neural networks (RNNs) (Salinas et al., 2020; Shen et al., 2020; Hochreiter & Schmidhuber, 1997) to transformers (Zhou et al., 2023; Liu et al., 2023; Nie et al., 2022; Zhang & Yan, 2022; Woo et al., 2022; Zhou et al., 2022; Liu et al., 2022b; Wu et al., 2022b; Xu et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2021; Kitaev et al., 2020; Li et al., 2019). Many models are hybrid solutions that blend aforementioned approaches.
Most of these methods intake time and then perform various combinations of normalization (Kim et al., 2021), frequency transformations (Wu et al., 2022a; Zhou et al., 2022), and patchification either along the time dimension (Liu et al., 2023; Zhang & Yan, 2022; Nie et al., 2022), or sensor dimension (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021). [^3] In time series analysis, sensors, channels, and variates are synonymous terms; in this paper we adopt the sensor terminology. Patch lengths range from a single time-step / sensor, also known as point-wise, to the length of the entire time series / all sensors. Time and sensor patch dependencies are then learned, via an attention mechanism, convolution, recurrence, or linear layer, across the temporal dimension, sensor dimension, or both the temporal and sensor dimensions (Zhang & Yan, 2022). For multisensor modeling, one can model all sensors jointly or independently (i.e., forecast each sensor independently) (Nie et al., 2022). These methods learn the underlying data representations end-to-end with the downstream task (e.g., forecasting).
Specialist-training, where models are only trained on a single time series domain, is the most common regime amongst prior work (Zhou et al., 2023; Wu et al., 2022a; Nie et al., 2022; Zhang & Yan, 2022). These specialist models are primarily evaluated via in-domain-testing, where the test set is from the same domain as the train set. Recently, some methods (Zhou et al., 2023; Liu et al., 2023) have begun to explore specialist zero-shot forecasting capabilities.
The time series analysis field is undergoing unification along both the modeling axis (Zhou et al., 2023; Wu et al., 2022a) and data representation axis (Franceschi et al., 2019; Tonekaboni et al., 2021; Yang & Hong, 2022; Barnum et al., 2020; Yue et al., 2022). Unified data representations, both statistical and learnt, have been more extensively studied in language and vision modeling (Gage, 1994; Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022). The vision modeling field distinguishes between discrete, learnt, tokens (Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022) and patches (Dosovitskiy et al., 2020). Patches have been studied in time series modeling (Zhou et al., 2023; Nie et al., 2022; Zhang & Yan, 2022). In this work, we propose to use discrete, learnt tokenized representations, which we show lead to strong performance in both specialist and generalist settings, as well as in-domain and zero-shot testing regimes.
3 Method
Our proposed discrete time series tokenization enables the design of general models across a variety of time series domains, tasks, and evaluation schemas, Figure 1. We design a single tokenizer architecture that is generally applicable without extensive data engineering while being suitable for varying data dimensionalities across different tasks. There are many possibilities for how to introduce a discrete time series tokenizer, we extensively study one such methodology that satisfies the aforementioned design criteria.
Data Engineering. Prior work in time series analysis leverages data engineering such as the use of auxiliary features (e.g. day of the month, or minute in the hour, etc.) (Chen et al., 2023; Salinas et al., 2020), or frequency transformations (Wu et al., 2022a; Zhou et al., 2022). We forego any data engineering and operate directly on time steps. This enables generalist-training as differing data domains have widely varying sampling rates leading to distinct auxiliary features and frequency profiles.
Varying Dimensionality. A time series dataset consists of $E$ examples (i.e. number of distinct recordings), $S$ sensor channels, and $T$ time steps, and can be formally expressed as $\{\mathbf{x}_{j}\}_{j=1}^{E}\subset\mathbb{R}^{S\times T}$. Even within a single task and single data domain where $S$ does not change, $E$ and $T$ take on a wide range of values. As an example, canonical forecasting predictions lengths range from $96$ to $720$ time steps. When moving to generalist-training, datasets additionally have wide ranging sensor dimensionalities $S$. Our tokenizer handles varying dimensionality across $E$, $S$, and $T$ by creating non-overlapping tokens along the time-dimension that are smaller than the dimension $T$.
Differing Tasks. There are numerous tasks to tackle in time series analysis. Three significant ones are imputation, anomaly detection, and forecasting. In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, and then reconstruct and impute $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$. In anomaly detection, models intake a corrupted time series $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and reconstruct the data $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$. The amount of corruption is considered known, at A%. In forecasting, models intake a time series $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict future readings $\mathbf{y}\in\mathbb{R}^{S\times T_{\textrm{out}}}$, where $S$ is the number of sensors and $T_{\textrm{in}},T_{\textrm{out}}$ signify the durations of the preceding and succeeding time series, respectively. Our tokenizer is performant across all tasks despite their distinct representational requirements.
TOTEM Implementation. To realize a single tokenizer architecture that enables generalist modeling across differing domains and tasks we take inspiration from the VQVAE (Van Den Oord et al., 2017). The original VQVAE leverages a dilated convolutional architecture with a stride of 2 and window-size of 4, similar to the WaveNet (Oord et al., 2016) dilated, causal, convolutional decoder. A dilated convolution skips inputs allowing a filter to operate on a larger input area / coarser scale. Utilizing dilated convolutions is an architectural decision rooted in the high sampling rates of raw audio waveforms (Oord et al., 2016; Van Den Oord et al., 2017). High sampling rates are not a trait shared by many time series domains.
When adapting the VQVAE for general time series analysis, the TOTEM VQVAE:
  1. 1.
    Operates directly on time steps; no data engineering.
  2. 2.
    Creates discrete, non-overlapping tokens along the time dimension of length $F$, where $F<T$, thereby promoting training and testing on variable length examples, $E$, sensors, $S$, and time steps $T$.
  3. 3.
    Maintains the same architecture and objective regardless of the downstream task.
  4. 4.
    Aims to capture maximal information within a large receptive field by: (1) using a strided non-causal convolutional architecture with no dilation, (2) training on long time series inputs, (3) pre-striding the data by a stride of 1 so the tokenizer learns from maximal inputs.
The TOTEM VQVAE consists of an encoder, quantizer, latent codebook, and decoder. It takes in a univariate time series $\{\mathbf{x}_{i}\in\mathbb{R}^{T}\}_{i=1}^{E\cdot S}$ obtained by flattening the sensor channel of the multivariate data. This makes TOTEM’s VQVAE sensor-agnostic, enabling TOTEM’s generalist-training and zero-shot-testing. The encoder $\mathcal{E}$ consists of strided 1D convolutions compressing the time series by a cumulative stride of $F$. $\mathcal{E}$ maps a univariate time series $\mathbf{x}\in\mathbb{R}^{T}$ to a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{\nicefrac{{T}}{{F}}\times D}$, where $D$ is the the hidden dimension. The latent codebook $\mathcal{C}=\{\mathbf{c}_{i}\}_{i=1}^{K}$ consists of $K$ $D$-dim codewords $\mathbf{c}_{i}\in\mathbb{R}^{D}$. During quantization, the codebook is used to replace $\mathbf{z}$ with $\hat{\mathbf{z}}\in\mathbb{R}^{\nicefrac{{T}}{{F}}\times D}$ such that $\hat{\mathbf{z}}_{j}=\mathbf{c}_{k}$, where $k=\operatorname*{arg\,min}_{i}||\mathbf{z}_{j}-c_{i}||_{2}$. The decoder $\mathcal{D}$ follows the reverse architecture of the encoder $\mathcal{E}$, consisting of 1D transpose convolutions with a cumulative stride of $\nicefrac{{1}}{{F}}$ mapping the quantized $\hat{\mathbf{z}}$ to a reconstructed time series $\hat{\mathbf{x}}=\mathcal{D}(\hat{\mathbf{z}})\in\mathbb{R}^{T}$. We learn $\mathcal{E}$, $\mathcal{D}$, and $\mathcal{C}$ by optimizing the objective $\mathcal{L}=\mathcal{L}_{\textrm{rec}}+\mathcal{L}_{\textrm{cmt}}$ consisting of a reconstruction loss $\mathcal{L}_{\textrm{rec}}~{}=~{}\frac{1}{E\cdot S}\sum_{i}||\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}||^{2}_{2}$ and a commitment loss $\mathcal{L}_{\textrm{cmt}}$, which allows the codebook to update despite the the non-differentiable $\operatorname*{arg\,min}$ operation during quantization. The final objective is $\mathcal{L}=\mathcal{L}_{\textrm{rec}}+\alpha\cdot\mathcal{L}_{\textrm{cmt}}$, where $\alpha$ is a scalar that weights the two losses. This objective does not change even when the underlying task, time series length, data masking, normalization schema, or data domain changes.
Notably imputation and anomaly detection can be directly solved with just TOTEM’s VQVAE, as they are fundamentally data representation tasks, whereas in forecasting further modeling is required, see Figure 2. In forecasting, the trained, frozen, codebook representation converts a sensor’s observed measurements $\mathbf{x}_{s}\in\mathbb{R}^{T_{\text{in}}}$ to a sequence of $\nicefrac{{T_{\text{in}}}}{{F}}$ discrete tokens. The forecaster transformer encoder processes these tokenized time series independently for each sensor, adding time-based positional encodings to each token along the time dimension. Using a series of multi-head attention layers, the model predicts the forecasted measurements $\bar{\mathbf{y}}_{s}\in\mathbb{R}^{T_{\textrm{out}}}$ for $s=1,...,S$, applying the attention mechanism along the time dimension $T$. In parallel, the forecaster takes in $\mathbf{x}_{s}$ and predicts the future’s mean, $\mu_{s}$, and standard deviation, $\sigma_{s}$, for each sensor $s=1,...,S$ to unnormalize the data. The final forecasted prediction is $\mathbf{y}_{s}=\sigma_{s}\cdot\bar{\mathbf{y}}_{s}+\mu_{s}$. The forecaster is trained in a supervised fashion by minimizing three smooth L1 losses between predictions $\{\bar{\mathbf{y}}_{s},\mu_{s},\sigma_{s}\}$ and their ground truth respectively.
Table 1: Specialist Imputation ()normal-↓($\downarrow$)( ↓ ). Across all datasets, metrics, and masking percentages, TOTEM has the highest AvgWins (52.1%), followed by GPT2 (35.4%). TOTEM values are means from 3 seeds; baseline values are from (Zhou et al., 2023; Wu et al., 2022a).
Model
TOTEM
GPT2
TiNet
Patch
ETS
FED
Stat
Auto
Inf
Re
LiTS
Dlin
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
W
12.5%
0.028
0.046
0.026
0.049
0.025
0.045
0.029
0.049
0.057
0.141
0.041
0.107
0.027
0.051
0.026
0.047
0.037
0.093
0.031
0.076
0.047
0.101
0.039
0.084
25%
0.029
0.047
0.028
0.052
0.029
0.052
0.031
0.053
0.065
0.155
0.064
0.163
0.029
0.056
0.030
0.054
0.042
0.100
0.035
0.082
0.052
0.111
0.048
0.103
37.5%
0.031
0.048
0.033
0.060
0.031
0.057
0.035
0.058
0.081
0.180
0.107
0.229
0.033
0.062
0.032
0.060
0.049
0.111
0.040
0.091
0.058
0.121
0.057
0.117
50%
0.033
0.052
0.037
0.065
0.034
0.062
0.038
0.063
0.102
0.207
0.183
0.312
0.037
0.068
0.037
0.067
0.053
0.114
0.046
0.099
0.065
0.133
0.066
0.134
E
12.5%
0.054
0.154
0.080
0.194
0.085
0.202
0.055
0.160
0.196
0.321
0.107
0.237
0.093
0.210
0.089
0.210
0.218
0.326
0.190
0.308
0.102
0.229
0.092
0.214
25%
0.059
0.160
0.087
0.203
0.089
0.206
0.065
0.175
0.207
0.332
0.120
0.251
0.097
0.214
0.096
0.220
0.219
0.326
0.197
0.312
0.121
0.252
0.118
0.247
37.5%
0.067
0.169
0.094
0.211
0.094
0.213
0.076
0.189
0.219
0.344
0.136
0.266
0.102
0.220
0.104
0.229
0.222
0.328
0.203
0.315
0.141
0.273
0.144
0.276
50%
0.079
0.183
0.101
0.220
0.100
0.221
0.091
0.208
0.235
0.357
0.158
0.284
0.108
0.228
0.113
0.239
0.228
0.331
0.210
0.319
0.160
0.293
0.175
0.305
m1
12.5%
0.049
0.125
0.017
0.085
0.019
0.092
0.041
0.130
0.067
0.188
0.035
0.135
0.026
0.107
0.034
0.124
0.047
0.155
0.032
0.126
0.075
0.180
0.058
0.162
25%
0.052
0.128
0.022
0.096
0.023
0.101
0.044
0.135
0.096
0.229
0.052
0.166
0.032
0.119
0.046
0.144
0.063
0.180
0.042
0.146
0.093
0.206
0.080
0.193
37.5%
0.055
0.132
0.029
0.111
0.029
0.111
0.049
0.143
0.133
0.271
0.069
0.191
0.039
0.131
0.057
0.161
0.079
0.200
0.063
0.182
0.113
0.231
0.103
0.219
50%
0.061
0.139
0.040
0.128
0.036
0.124
0.055
0.151
0.186
0.323
0.089
0.218
0.047
0.145
0.067
0.174
0.093
0.218
0.082
0.208
0.134
0.255
0.132
0.248
m2
12.5%
0.016
0.078
0.017
0.076
0.018
0.080
0.026
0.094
0.108
0.239
0.056
0.159
0.021
0.088
0.023
0.092
0.133
0.270
0.108
0.228
0.034
0.127
0.062
0.166
25%
0.017
0.081
0.020
0.080
0.020
0.085
0.028
0.099
0.164
0.294
0.080
0.195
0.024
0.096
0.026
0.101
0.135
0.272
0.136
0.262
0.042
0.143
0.085
0.196
37.5%
0.018
0.084
0.022
0.087
0.023
0.091
0.030
0.104
0.237
0.356
0.110
0.231
0.027
0.103
0.030
0.108
0.155
0.293
0.175
0.300
0.051
0.159
0.106
0.222
50%
0.020
0.088
0.025
0.095
0.026
0.098
0.034
0.110
0.323
0.421
0.156
0.276
0.030
0.108
0.035
0.119
0.200
0.333
0.211
0.329
0.059
0.174
0.131
0.247
h1
12.5%
0.119
0.212
0.043
0.140
0.057
0.159
0.093
0.201
0.126
0.263
0.070
0.190
0.060
0.165
0.074
0.182
0.114
0.234
0.074
0.194
0.240
0.345
0.151
0.267
25%
0.127
0.220
0.054
0.156
0.069
0.178
0.107
0.217
0.169
0.304
0.106
0.236
0.080
0.189
0.090
0.203
0.140
0.262
0.102
0.227
0.265
0.364
0.180
0.292
37.5%
0.138
0.230
0.072
0.180
0.084
0.196
0.120
0.230
0.220
0.347
0.124
0.258
0.102
0.212
0.109
0.222
0.174
0.293
0.135
0.261
0.296
0.382
0.215
0.318
50%
0.157
0.247
0.107
0.216
0.102
0.215
0.141
0.248
0.293
0.402
0.165
0.299
0.133
0.240
0.137
0.248
0.215
0.325
0.179
0.298
0.334
0.404
0.257
0.347
h2
12.5%
0.040
0.129
0.039
0.125
0.040
0.130
0.057
0.152
0.187
0.319
0.095
0.212
0.042
0.133
0.044
0.138
0.305
0.431
0.163
0.289
0.101
0.231
0.100
0.216
25%
0.041
0.131
0.044
0.135
0.046
0.141
0.061
0.158
0.279
0.390
0.137
0.258
0.049
0.147
0.050
0.149
0.322
0.444
0.206
0.331
0.115
0.246
0.127
0.247
37.5%
0.043
0.136
0.051
0.147
0.052
0.151
0.067
0.166
0.400
0.465
0.187
0.304
0.056
0.158
0.060
0.163
0.353
0.462
0.252
0.370
0.126
0.257
0.158
0.276
50%
0.047
0.142
0.059
0.158
0.060
0.162
0.073
0.174
0.602
0.572
0.232
0.341
0.065
0.170
0.068
0.173
0.369
0.472
0.316
0.419
0.136
0.268
0.183
0.299
AvgWins
52.1% 35.4% 18.8% 0% 0% 0% 0% 0% 0% 0% 0% 0%
Table 2: Generalist Imputation ()normal-↓($\downarrow$)( ↓ ). TOTEM & GPT2 simultaneously train on all in domain datasets, 3 seeds each. A. In-Domain Performance. TOTEM has the highest AvgWins at 58.3%. B. Zero-Shot Performance. We test on unseen datasets zero-shot. TOTEM again has the highest AvgWins at 80.0%.
[t] A. In-Domain Performance Model TOTEM GPT2 Metric MSE MAE MSE MAE W 12.5% 0.029 0.060 0.029 0.045 25% 0.030 0.060 0.033 0.048 37.5% 0.032 0.062 0.037 0.054 50% 0.036 0.067 0.043 0.061 E 12.5% 0.065 0.171 0.080 0.186 25% 0.071 0.179 0.091 0.197 37.5% 0.080 0.189 0.108 0.213 50% 0.095 0.205 0.132 0.236 m1 12.5% 0.041 0.132 0.052 0.141 25% 0.044 0.135 0.065 0.154 37.5% 0.048 0.139 0.085 0.171 50% 0.058 0.152 0.117 0.196 m2 12.5% 0.040 0.125 0.029 0.095 25% 0.041 0.126 0.033 0.101 37.5% 0.043 0.129 0.038 0.110 50% 0.048 0.136 0.045 0.121 h1 12.5% 0.100 0.201 0.113 0.217 25% 0.108 0.209 0.131 0.231 37.5% 0.122 0.220 0.153 0.247 50% 0.144 0.237 0.182 0.266 h2 12.5% 0.075 0.175 0.067 0.155 25% 0.076 0.177 0.071 0.160 37.5% 0.093 0.195 0.077 0.167 50% 0.089 0.192 0.086 0.179 AvgWins 58.3% 43.8%
B. Zero-Shot Performance
Model
TOTEM
GPT2
Metric
MSE
MAE
MSE
MAE
N2
12.5%
0.029
0.120
0.047
0.145
25%
0.033
0.127
0.064
0.164
37.5%
0.041
0.139
0.090
0.191
50%
0.056
0.160
0.131
0.228
N5
12.5%
0.017
0.085
0.021
0.095
25%
0.019
0.090
0.028
0.107
37.5%
0.022
0.098
0.039
0.123
50%
0.029
0.110
0.055
0.145
R
12.5%
0.071
0.109
0.093
0.119
25%
0.087
0.117
0.125
0.134
37.5%
0.112
0.129
0.167
0.154
50%
0.148
0.147
0.220
0.182
B
12.5%
0.632
0.642
0.392
0.496
25%
0.693
0.665
0.444
0.523
37.5%
0.761
0.692
0.498
0.553
50%
0.827
0.718
0.591
0.599
S
12.5%
0.057
0.160
0.070
0.173
25%
0.061
0.168
0.084
0.189
37.5%
0.069
0.178
0.103
0.209
50%
0.082
0.193
0.128
0.234
AvgWins
80.0% 20.0%
4 Experimental Setup
Through experiments in imputation (§5), anomaly detection (§6), and forecasting (§7), our goal is to explore the efficacy of TOTEM on standard benchmark datasets and tasks, and domain general settings. To briefly refresh: specialist refers to training on a single domain (Tables 1, 3, 5). Generalist refers to training on multiple domains (Tables 2, 4, 6). Finally, in-domain refers to testing on the training domain, and zero-shot to testing on a separate domain from training.
Figure 2: Forecaster Modeling. The forecasting task requires modeling beyond the VQVAE. We leverage TOTEM’s pretrained, learnt, discrete codes as a the input data representation and train a transformer encoder. We add positional embeddings along the time dimension, and use linear layers before the final output as well as to un-normalize the resulting forecast.
For all experiments & models, we run three seeds and report the mean; standard deviations are reported in the Appendix A. Following the field standard, we bold the best metric in all tables. Evaluation metrics differ across tasks. We report mean squared error MSE ($\downarrow$), mean absolute error MAE ($\downarrow$), precision P ($\uparrow$), recall R ($\uparrow$), and F1 score ($\uparrow$); ($\downarrow$) means lower is better, ($\uparrow$) means higher performance is better. Given the varied metrics we calculate the average number of best results, or AvgWins , for each method and highlight the best, second best, and third best methods.
Baselines. We compare to two families of approaches: methods designed for multiple tasks (multitask) – TOTEM belongs in this category – and methods designed for a specific task (singletask), and are adapted to other tasks.
We compare against two recent multitask models, the transformer based GPT2 (Zhou et al., 2023) and the convolutional TimesNet [TiNet] (Wu et al., 2022a). For singletask models we compare against PatchTST [Patch] (Nie et al., 2022), ETSFormer [ETS] (Woo et al., 2022), Fedformer [FED] (Zhou et al., 2022), Non-stationary trans. [Stat] (Liu et al., 2022b), Autoformer [Auto] (Wu et al., 2021), Informer [Inf] (Zhou et al., 2021), Reformer [Re] (Kitaev et al., 2020), LightTS [LiTS] (Zhang et al., 2022), DLinear [DLin] (Zeng et al., 2023), Anomaly trans. [ATran](Xu et al., 2021), Pyraformer [Pyra] (Liu et al., 2021), LogTrans. [LogTr] (Li et al., 2019), Trans. [Trans] (Vaswani et al., 2017), Crossformer [Cross] (Zhang & Yan, 2022), TiDE (Das et al., 2023), RLinear [RLin] (Li et al., 2023), SciNet [SCi] (Liu et al., 2022a), & iTrans. [iTrans] (Liu et al., 2023).
Datasets. We leverage 12 benchmark datasets: weather [W], electricity [E], traffic [T], ETTm1 [m1], ETTm2 [m2], ETTh1 [h1], ETTh2 [h2], SMD, MSL, SMAP, SWAT, PSM that are commonly used for imputation, anomaly detection and forecasting (Zhou et al., 2023; Wu et al., 2022a; Xu et al., 2021; Zhang & Yan, 2022; Nie et al., 2022). For the zero shot settings, we leverage 5 benchmark datasets: neuro2 [N2], neuro5 [N5] (Peterson et al., 2022), and saugeen river flow [R], U.S. births [B], and sunspot [S] (Godahewa et al., 2021). 17 datasets in total.
5 Imputation
In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, and then reconstruct and impute $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ We experiment with four canonical masking percentages at $12.5\%,25\%,37.5\%,50\%$, and report MSE and MAE .
Specialist. In Table 1 we compare TOTEM to baselines. All models are trained and evaluated on the same dataset (in-domain). TOTEM has the highest AvgWins with 52.1%, followed by GPT2 at 35.4%, and TiNet at 18.8%. TOTEM performance for m1 and h1 is lower; notably these datasets are the minute and hour resampling of the same raw data respectively. We investigate and discuss TOTEM’s success across different domains in Table 9.
Generalist. In Table 2 we compare TOTEM to GPT2 (best performing models above), when both models are trained on the aggregate of W, E, m1, m2, h1, h2. We test them on the in-domain and zero-shot test sets. TOTEM outperforms GPT2 in-domain, 58.3% vs. 43.8% , and by a much larger margin in zero-shot, 80% vs. 20%.
TOTEM’s performance across all experiments demonstrate that tokens are a performant representation for imputation.
Table 3: Specialist Anomaly Detection ()normal-↑($\uparrow$)( ↑ ). TOTEM has the highest AvgWins at 33.3% followed by a five-way tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%. Some prior methods use the test set as a validation set for early stopping of the learning algorithm, which can inflate performance. We do not adopt this practice and train TOTEM for a set number of iterations.
Model
TOTEM
GPT2
TiNet
ATran
Patch
ETS
FED
Stat
Auto
Pyra
Inf
Re
LogTr
Trans
LiTS
DLin
F1
SMD
79.62
86.89
84.61
85.49
84.62
83.13
85.08
84.62
85.11
83.04
81.65
75.32
76.21
79.56
82.53
77.10
MSL
82.58
82.45
81.84
83.31
78.70
85.03
78.57
77.50
79.05
84.86
84.06
84.40
79.57
78.68
78.95
84.88
SMAP
94.02
72.88
69.39
71.18
68.82
69.50
70.76
71.09
71.12
71.09
69.92
70.40
69.97
69.70
69.21
69.26
SWAT
94.27
94.23
93.02
83.10
85.72
84.91
93.19
79.88
92.74
91.78
81.43
82.80
80.52
80.37
93.33
87.52
PSM
95.87
97.13
97.34
79.40
96.08
91.76
97.23
97.29
93.29
82.08
77.10
73.61
76.74
76.07
97.15
93.55
R
SMD
76.06
84.98
81.54
82.23
82.14
79.23
82.39
81.21
82.35
80.61
77.23
69.24
70.13
76.13
78.42
71.52
MSL
82.85
82.91
75.36
87.37
70.96
84.93
80.07
89.14
80.92
85.93
86.48
83.31
87.37
87.37
75.78
85.42
SMAP
94.04
60.95
56.40
58.11
55.46
55.75
58.10
59.02
58.62
57.71
57.13
57.44
57.59
57.12
55.27
55.41
SWAT
95.91
96.34
95.40
97.32
80.94
80.36
96.42
96.75
95.81
96.00
96.75
96.53
97.32
96.53
94.72
95.30
PSM
94.21
95.68
96.20
94.72
93.47
85.28
97.16
96.76
88.15
96.02
96.33
95.38
98.00
96.56
95.97
89.26
P
SMD
83.54
88.89
87.91
88.91
87.26
87.44
87.95
88.33
88.06
85.61
86.60
82.58
83.46
83.58
87.10
83.62
MSL
82.32
82.00
89.54
79.61
88.34
85.13
77.14
68.55
77.27
83.81
81.77
85.51
73.05
71.57
82.40
84.34
SMAP
94.00
90.60
90.14
91.85
90.64
92.25
90.47
89.37
90.40
92.54
90.11
90.91
89.15
89.37
92.58
92.32
SWAT
92.68
92.20
90.75
72.51
91.10
90.02
90.17
68.03
89.85
87.92
70.29
72.50
68.67
68.84
91.98
80.91
PSM
97.58
98.62
98.51
68.35
98.84
99.31
97.31
97.82
99.08
71.67
64.27
59.93
63.06
62.75
98.37
98.28
AvgWins
33.3% 13.3% 13.3% 13.3% 0% 13.3% 0% 6.7% 0% 0% 0% 0% 13.3% 0% 0% 0%
Table 4: Generalist Anomaly Detection ()normal-↑($\uparrow$)( ↑ ). We train TOTEM & GPT2 on all datasets and then perform in-domain and zero-shot evaluations. A. In-Domain Performance. TOTEM outperforms GPT2: 80.0% vs. 20.0%. B. Zero-Shot Performance. TOTEM again outperforms GPT2: 73.3% vs. 26.7%.
A. In-Domain Performance
Model
TOTEM
GPT2
F1
SMD
78.64
79.73
MSL
83.29
80.17
SMAP
92.51
67.05
SWAT
94.37
89.62
PSM
95.78
90.47
R
SMD
72.07
73.42
MSL
82.96
78.48
SMAP
91.48
53.42
SWAT
96.13
87.53
PSM
93.90
87.76
P
SMD
86.66
87.44
MSL
83.64
81.95
SMAP
93.56
90.01
SWAT
92.68
91.83
PSM
97.74
93.39
AvgWins
80.0% 20.0%
B. Zero-Shot Performance
Model
TOTEM
GPT2
F1
N2
51.29
39.02
N5
51.28
42.19
R
49.39
36.14
B
49.15
20.81
S
52.17
38.12
R
N2
76.88
33.69
N5
76.84
36.77
R
70.49
29.66
B
73.71
17.67
S
77.36
31.83
P
N2
38.49
46.43
N5
38.48
49.58
R
38.02
46.30
B
36.86
25.33
S
39.35
47.72
AvgWins
73.3% 26.7%
6 Anomaly Detection
In anomaly detection, models intake a corrupted time series $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and reconstruct the data $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, where the amount of corruption is considered known, at A%. We report % Precision P ($\uparrow$), Recall R ($\uparrow$), and F1 Score ($\uparrow$).
The standard practice in machine learning, which we adopt, is to have a held out test set that is not used for tuning the model or learning algorithm. One aspect that makes comparing with several prior works challenging is that they use the test set as a validation set for early stopping of the learning algorithm, which can often inflate their performance. Despite this inconsistency, we compare our performance against these reported performances, whenever available.
Specialist. In Table 3 we evaluate TOTEM against numerous specialist baselines. TOTEM has the highest AvgWins at 33.3% followed by a five-way tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%.
Generalist. In Table 4 we compare generalist-trained TOTEM and GPT2. On the in-domain test sets TOTEM outperforms GPT2: 80% vs. 20%. In the zero-shot test sets TOTEM outperforms GPT2: 73.3% vs. 26.7%.
TOTEM’s AvgWins across the specialist and generalist settings demonstrate that tokens are a performant representation for anomaly detection.
Table 5: Specialist Forecasting ()normal-↓($\downarrow$)( ↓ ). TOTEM has the best AvgWins (28.6%), followed by iTrans (26.8%). Notably, TOTEM has first place finishes in 5 datasets, while iTrans’ first places are concentrated in only electricity and traffic. All models have lookback $T_{in}=96$.
Model
TOTEM
GPT2
TiNet
iTrans
Patch
Cross
FED
Stat
TiDE
RLin
DLin
SCi
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
W
96
0.165
0.208
0.184
0.224
0.172
0.220
0.174
0.214
0.177
0.218
0.158
0.230
0.217
0.296
0.173
0.223
0.202
0.261
0.192
0.232
0.196
0.255
0.221
0.306
192
0.207
0.250
0.231
0.263
0.219
0.261
0.221
0.254
0.225
0.259
0.206
0.277
0.276
0.336
0.245
0.285
0.242
0.298
0.240
0.271
0.237
0.296
0.261
0.340
336
0.257
0.291
0.285
0.302
0.280
0.306
0.278
0.296
0.278
0.297
0.272
0.335
0.339
0.380
0.321
0.338
0.287
0.335
0.292
0.307
0.283
0.335
0.309
0.378
720
0.326
0.340
0.362
0.351
0.365
0.359
0.358
0.349
0.354
0.348
0.398
0.418
0.403
0.428
0.414
0.410
0.351
0.386
0.364
0.353
0.345
0.381
0.377
0.427
E
96
0.178
0.263
0.186
0.272
0.168
0.272
0.148
0.240
0.195
0.285
0.219
0.314
0.193
0.308
0.169
0.273
0.237
0.329
0.201
0.281
0.197
0.282
0.247
0.345
192
0.187
0.272
0.190
0.278
0.184
0.289
0.162
0.253
0.199
0.289
0.231
0.322
0.201
0.315
0.182
0.286
0.236
0.330
0.201
0.283
0.196
0.285
0.257
0.355
336
0.199
0.285
0.204
0.291
0.198
0.300
0.178
0.269
0.215
0.305
0.246
0.337
0.214
0.329
0.200
0.304
0.249
0.344
0.215
0.298
0.209
0.301
0.269
0.369
720
0.236
0.318
0.245
0.324
0.220
0.320
0.225
0.317
0.256
0.337
0.280
0.363
0.246
0.355
0.222
0.321
0.284
0.373
0.257
0.331
0.245
0.333
0.299
0.390
T
96
0.523
0.303
0.471
0.311
0.593
0.321
0.395
0.268
0.544
0.359
0.522
0.290
0.587
0.366
0.612
0.338
0.805
0.493
0.649
0.389
0.650
0.396
0.788
0.499
192
0.530
0.303
0.479
0.312
0.617
0.336
0.417
0.276
0.540
0.354
0.530
0.293
0.604
0.373
0.613
0.340
0.756
0.474
0.601
0.366
0.598
0.370
0.789
0.505
336
0.549
0.311
0.490
0.317
0.629
0.336
0.433
0.283
0.551
0.358
0.558
0.305
0.621
0.383
0.618
0.328
0.762
0.477
0.609
0.369
0.605
0.373
0.797
0.508
720
0.598
0.331
0.524
0.336
0.640
0.350
0.467
0.302
0.586
0.375
0.589
0.328
0.626
0.382
0.653
0.355
0.719
0.449
0.647
0.387
0.645
0.394
0.841
0.523
m1
96
0.320
0.347
0.328
0.363
0.338
0.375
0.334
0.368
0.329
0.367
0.404
0.426
0.379
0.419
0.386
0.398
0.364
0.387
0.355
0.376
0.345
0.372
0.418
0.438
192
0.379
0.382
0.368
0.382
0.374
0.387
0.377
0.391
0.367
0.385
0.450
0.451
0.426
0.441
0.459
0.444
0.398
0.404
0.391
0.392
0.380
0.389
0.439
0.450
336
0.406
0.402
0.400
0.404
0.410
0.411
0.426
0.420
0.399
0.410
0.532
0.515
0.445
0.459
0.495
0.464
0.428
0.425
0.424
0.415
0.413
0.413
0.490
0.485
720
0.471
0.438
0.462
0.440
0.478
0.450
0.491
0.459
0.454
0.439
0.666
0.589
0.543
0.490
0.585
0.516
0.487
0.461
0.487
0.450
0.474
0.453
0.595
0.550
m2
96
0.176
0.253
0.178
0.263
0.187
0.267
0.180
0.264
0.175
0.259
0.287
0.366
0.203
0.287
0.192
0.274
0.207
0.305
0.182
0.265
0.193
0.292
0.286
0.377
192
0.247
0.302
0.245
0.307
0.249
0.309
0.250
0.309
0.241
0.302
0.414
0.492
0.269
0.328
0.280
0.339
0.290
0.364
0.246
0.304
0.284
0.362
0.399
0.445
336
0.317
0.348
0.307
0.346
0.321
0.351
0.311
0.348
0.305
0.343
0.597
0.542
0.325
0.366
0.334
0.361
0.377
0.422
0.307
0.342
0.369
0.427
0.637
0.591
720
0.426
0.410
0.410
0.409
0.408
0.403
0.412
0.407
0.402
0.400
1.730
1.042
0.421
0.415
0.417
0.413
0.558
0.524
0.407
0.398
0.554
0.522
0.960
0.735
h1
96
0.380
0.394
0.379
0.397
0.384
0.402
0.386
0.405
0.414
0.419
0.423
0.448
0.376
0.419
0.513
0.491
0.479
0.464
0.386
0.395
0.386
0.400
0.654
0.599
192
0.434
0.427
0.438
0.427
0.436
0.429
0.441
0.436
0.460
0.445
0.471
0.474
0.420
0.448
0.534
0.504
0.525
0.492
0.437
0.424
0.437
0.432
0.719
0.631
336
0.490
0.459
0.474
0.448
0.491
0.469
0.487
0.458
0.501
0.466
0.570
0.546
0.459
0.465
0.588
0.535
0.565
0.515
0.479
0.446
0.481
0.459
0.778
0.659
720
0.539
0.513
0.496
0.475
0.521
0.500
0.503
0.491
0.500
0.488
0.653
0.621
0.506
0.507
0.643
0.616
0.594
0.558
0.481
0.470
0.519
0.516
0.836
0.699
h2
96
0.293
0.338
0.295
0.348
0.340
0.374
0.297
0.349
0.302
0.348
0.745
0.584
0.358
0.397
0.476
0.458
0.400
0.440
0.288
0.338
0.333
0.387
0.707
0.621
192
0.375
0.390
0.384
0.402
0.402
0.414
0.380
0.400
0.388
0.400
0.877
0.656
0.429
0.439
0.512
0.493
0.528
0.509
0.374
0.390
0.477
0.476
0.860
0.689
336
0.422
0.431
0.418
0.432
0.452
0.452
0.428
0.432
0.426
0.433
1.043
0.731
0.496
0.487
0.552
0.551
0.643
0.571
0.415
0.426
0.594
0.541
1.000
0.744
720
0.610
0.567
0.423
0.446
0.462
0.468
0.427
0.445
0.431
0.446
1.104
0.763
0.463
0.474
0.562
0.560
0.874
0.679
0.420
0.440
0.831
0.657
1.249
0.838
AvgWins
28.6% 1.8% 1.8% 26.8% 14.3% 3.6% 5.4% 0% 0% 25% 0% 0%
Table 6: Generalist Forecasting ()normal-↓($\downarrow$)( ↓ ). Here we evaluate the generalist TOTEM and GPT2 models. A. In-Domain Performance. TOTEM outperforms GPT2: 67.9% to 33.9%. B. Zero-Shot Performance. TOTEM outperforms GPT2: 90.0% to 12.5%.
A. In-Domain Performance
Model
TOTEM
GPT2
Metric
MSE
MAE
MSE
MAE
W
96
0.172
0.216
0.201
0.237
192
0.217
0.256
0.247
0.275
336
0.266
0.295
0.298
0.311
720
0.334
0.342
0.372
0.360
E
96
0.179
0.264
0.194
0.278
192
0.181
0.267
0.199
0.284
336
0.196
0.283
0.214
0.300
720
0.230
0.314
0.255
0.331
T
96
0.507
0.284
0.484
0.320
192
0.511
0.282
0.488
0.320
336
0.535
0.292
0.502
0.326
720
0.580
0.309
0.534
0.343
m1
96
0.374
0.384
0.487
0.468
192
0.400
0.399
0.516
0.480
336
0.432
0.424
0.548
0.499
720
0.487
0.460
0.581
0.511
m2
96
0.198
0.275
0.243
0.315
192
0.266
0.319
0.297
0.346
336
0.365
0.377
0.349
0.376
720
0.588
0.511
0.439
0.423
h1
96
0.382
0.404
0.421
0.408
192
0.463
0.435
0.480
0.436
336
0.507
0.463
0.518
0.453
720
0.517
0.500
0.517
0.467
h2
96
0.307
0.345
0.298
0.343
192
0.406
0.403
0.381
0.392
336
0.505
0.460
0.406
0.419
720
0.661
0.557
0.423
0.438
AvgWins
67.9% 33.9%
B. Zero-Shot Performance
Model
TOTEM
GPT2
Metric
MSE
MAE
MSE
MAE
N2
96
1.138
0.777
1.332
0.830
192
1.149
0.785
1.416
0.863
336
1.092
0.770
1.358
0.851
720
1.045
0.754
1.308
0.840
N5
96
0.483
0.484
0.528
0.499
192
0.495
0.491
0.578
0.524
336
0.468
0.483
0.548
0.515
720
0.451
0.477
0.537
0.511
R
96
1.120
0.582
1.465
0.725
192
1.242
0.635
1.638
0.785
336
1.237
0.626
1.601
0.769
720
1.182
0.604
1.552
0.760
B
96
0.805
0.739
0.838
0.762
192
0.836
0.752
0.837
0.752
336
0.809
0.748
0.792
0.738
720
0.896
0.794
0.927
0.806
S
96
0.446
0.482
0.443
0.478
192
0.462
0.491
0.481
0.499
336
0.521
0.525
0.541
0.533
720
0.717
0.625
0.773
0.643
AvgWins
90.0% 12.5%
7 Forecasting
In forecasting, models intake a time series $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict future readings $\mathbf{y}\in\mathbb{R}^{S\times T_{\textrm{out}}}$, where $S$ is the number of sensors and $T_{\textrm{in}},T_{\textrm{out}}$ signify the durations of the preceding and succeeding time series, respectively. The pairs $(\mathbf{x},\mathbf{y})$ are generated by striding the original time series data.
All models have a lookback of $T_{\textrm{in}}=96$, with prediction lengths $T_{\textrm{out}}=\{96,192,336,720\}$. Numbers for other methods are from (Liu et al., 2023). We run GPT2 with $T_{\textrm{in}}=96$ as they originally report varying, dataset-specific, lookback lengths. We report MSE ($\downarrow$) and MAE ($\downarrow$).
Specialist. From Table 5 we find that TOTEM achieves the highest AvgWins at 28.6% followed by iTrans at 26.8%. TOTEM has first finishes in five datasets while iTrans’ first finishes are concentrated in only electricity and traffic.
Generalist. In Table 6 we compare generalist TOTEM and GPT2. TOTEM outperforms GPT2 for both in-domain (67.9% vs. 33.9%) and zero-shot (90.0% vs. 12.5%).
TOTEM’s AvgWins forecasting performance across the training and testing regimes demonstrates that tokens are a performant representation for forecasting.
8 Ablations
Table 7: Ablations ()normal-↓($\downarrow$)( ↓ ). Across the Tokens vs. Time (TvT) experiments tokens out perform time. (A) specialist: 67.9% to 39.3%, (B) in-domain generalist: 78.6% to 23.2% , and (C) zero-shot generalist: 67.5% to 35%. (D) As the codebook size $K$ increases the VQVAE reconstruction performance improves.
A. TvT Specialist
Model
TOTEM
TimeTOTEM
Metric
MSE
MAE
MSE
MAE
W
96
0.165
0.208
0.164
0.209
192
0.207
0.250
0.209
0.251
336
0.257
0.291
0.261
0.293
720
0.326
0.340
0.332
0.340
E
96
0.178
0.263
0.179
0.262
192
0.187
0.272
0.185
0.269
336
0.199
0.285
0.204
0.289
720
0.236
0.318
0.244
0.325
T
96
0.523
0.303
0.528
0.310
192
0.530
0.303
0.500
0.349
336
0.549
0.311
0.531
0.365
720
0.598
0.331
0.578
0.398
m1
96
0.320
0.347
0.326
0.355
192
0.379
0.382
0.377
0.386
336
0.406
0.402
0.409
0.409
720
0.471
0.438
0.469
0.441
m2
96
0.176
0.253
0.176
0.254
192
0.247
0.302
0.247
0.303
336
0.317
0.348
0.318
0.350
720
0.426
0.410
0.419
0.411
h1
96
0.380
0.394
0.377
0.395
192
0.434
0.427
0.428
0.428
336
0.490
0.459
0.480
0.462
720
0.539
0.513
0.530
0.522
h2
96
0.293
0.338
0.294
0.338
192
0.375
0.390
0.373
0.389
336
0.422
0.431
0.423
0.433
720
0.610
0.567
0.591
0.556
AvgWins
67.9% 39.3%
B. TvT In-Domain Generalist
Model
TOTEM
TimeTOTEM
Metric
MSE
MAE
MSE
MAE
W
96
0.172
0.216
0.173
0.218
192
0.217
0.256
0.218
0.261
336
0.266
0.295
0.267
0.299
720
0.334
0.342
0.337
0.347
E
96
0.179
0.264
0.183
0.267
192
0.181
0.267
0.189
0.275
336
0.196
0.283
0.204
0.291
720
0.230
0.314
0.242
0.325
T
96
0.507
0.284
0.517
0.293
192
0.511
0.282
0.526
0.296
336
0.535
0.292
0.552
0.304
720
0.580
0.309
0.602
0.326
m1
96
0.374
0.384
0.428
0.420
192
0.400
0.399
0.438
0.427
336
0.432
0.424
0.469
0.447
720
0.487
0.460
0.546
0.493
m2
96
0.198
0.275
0.207
0.286
192
0.266
0.319
0.269
0.325
336
0.365
0.377
0.358
0.377
720
0.588
0.511
0.521
0.482
h1
96
0.382
0.404
0.401
0.410
192
0.463
0.435
0.453
0.441
336
0.507
0.463
0.496
0.468
720
0.517
0.500
0.518
0.510
h2
96
0.307
0.345
0.305
0.346
192
0.406
0.403
0.396
0.402
336
0.505
0.460
0.492
0.458
720
0.661
0.557
0.599
0.531
AvgWins
78.6% 23.2%
C. TvT Zero-Shot Generalist
Model
TOTEM
TimeTOTEM
Metric
MSE
MAE
MSE
MAE
N2
96
1.138
0.777
1.127
0.773
192
1.149
0.785
1.169
0.793
336
1.092
0.770
1.115
0.780
720
1.045
0.754
1.070
0.766
N5
96
0.483
0.484
0.481
0.483
192
0.495
0.491
0.508
0.500
336
0.468
0.483
0.481
0.491
720
0.451
0.477
0.467
0.488
R
96
1.120
0.582
1.102
0.578
192
1.242
0.635
1.207
0.628
336
1.237
0.626
1.190
0.613
720
1.182
0.604
1.149
0.596
B
96
0.805
0.739
0.825
0.751
192
0.836
0.752
0.847
0.761
336
0.809
0.748
0.831
0.764
720
0.896
0.794
0.928
0.813
S
96
0.446
0.482
0.446
0.481
192
0.462
0.491
0.478
0.499
336
0.521
0.525
0.535
0.532
720
0.717
0.625
0.736
0.631
AvgWins
67.5% 35.0%
D. Codebook Size Ablations
Codebook Size $K$
32
256
512
MSE
All
0.0451
0.0192
0.0184
T
0.0312
0.0120
0.0101
E
0.0463
0.0209
0.0152
W
0.0393
0.0161
0.0128
MAE
All
0.1460
0.0937
0.0913
T
0.1204
0.0749
0.0685
E
0.1520
0.1027
0.0878
W
0.1122
0.0673
0.0607
AvgWins
0%
0%
100%
Tokens vs. Time. To evaluate if tokens enable TOTEM’s performance, we implement TimeTOTEM. TimeTOTEM has the identical architecture to TOTEM, except we replace the VQVAE with an MLP trained end-to-end with the downstream forecaster. We compare Totem vs. TimeTOTEM in the specialist in-domain, and generalist in-domain and zero-shot regimes (Table 7). In all cases TOTEM outperforms TimeTOTEM - specialist: 67.9% vs. 39.3%, generalist in-domain: 78.6% vs. 23.2%, generalist zero-shot: 67.5% vs. 35.0%. TOTEM’s performance demonstrates that tokens, when compared to time, lead to better performance.
Codebook Size. In Table 7 we explore the affect of the codebook size, $K$, on the VQVAE’s MSE and MAE reconstruction performance. As expected, we find that as $K$ increases from 32 to 256 to 512 the reconstruction performance improves.
Table 8: Generalist codes beat specialist codes: 66.1% vs 57.1%.
Codebook
Specialist
Generalist
Generalist
Forecaster
Specialist
Specialist
Generalist
Metric
MSE
MAE
MSE
MAE
MSE
MAE
W
96
0.165
0.208
0.164
0.208
0.172
0.216
192
0.207
0.250
0.208
0.251
0.217
0.256
336
0.257
0.291
0.258
0.290
0.266
0.295
720
0.326
0.340
0.329
0.338
0.334
0.342
E
96
0.178
0.263
0.178
0.263
0.179
0.264
192
0.187
0.272
0.187
0.273
0.181
0.267
336
0.199
0.285
0.199
0.285
0.196
0.283
720
0.236
0.318
0.238
0.320
0.230
0.314
T
96
0.523
0.303
0.521
0.301
0.507
0.284
192
0.530
0.303
0.530
0.303
0.511
0.282
336
0.549
0.311
0.555
0.313
0.535
0.292
720
0.598
0.331
0.605
0.337
0.580
0.309
m1
96
0.320
0.347
0.328
0.352
0.374
0.384
192
0.379
0.382
0.377
0.383
0.400
0.399
336
0.406
0.402
0.408
0.404
0.432
0.424
720
0.471
0.438
0.470
0.440
0.487
0.460
m2
96
0.176
0.253
0.175
0.253
0.198
0.275
192
0.247
0.302
0.247
0.302
0.266
0.319
336
0.317
0.348
0.318
0.348
0.365
0.377
720
0.426
0.410
0.427
0.410
0.588
0.511
h1
96
0.380
0.394
0.382
0.395
0.382
0.404
192
0.434
0.427
0.437
0.427
0.463
0.435
336
0.490
0.459
0.490
0.460
0.507
0.463
720
0.539
0.513
0.536
0.512
0.517
0.500
h2
96
0.293
0.338
0.294
0.339
0.307
0.345
192
0.375
0.390
0.375
0.391
0.406
0.403
336
0.422
0.431
0.421
0.431
0.505
0.460
720
0.610
0.567
0.610
0.567
0.661
0.557
AvgWins
57.1% 66.1%
Table 9: Zero Shot Vignette: Training Size & Diversity
Model
TOTEM
TOTEM
TOTEM
Generalist
Specialist
Specialist
Train Domain
ALL
Traffic
Electricity
Sensor Num ($S$)
862
321
Raw Length ($T$)
17544
26304
Train Size
17.6M
10.2M
5.8M
Metric
MSE
MAE
MSE
MAE
MSE
MAE
N2
96
1.138
0.777
1.194
0.798
1.193
0.802
192
1.149
0.785
1.218
0.808
1.300
0.845
336
1.092
0.770
1.190
0.804
1.260
0.837
720
1.045
0.754
1.117
0.784
1.234
0.832
N5
96
0.483
0.484
0.515
0.505
0.489
0.490
192
0.495
0.491
0.535
0.514
0.555
0.527
336
0.468
0.483
0.524
0.513
0.538
0.525
720
0.451
0.477
0.500
0.507
0.533
0.527
R
96
1.120
0.582
1.171
0.635
1.141
0.579
192
1.242
0.635
1.273
0.673
1.297
0.652
336
1.237
0.626
1.232
0.653
1.247
0.628
720
1.182
0.604
1.198
0.642
1.236
0.633
B
96
0.805
0.739
0.812
0.749
0.820
0.756
192
0.836
0.752
0.858
0.767
0.843
0.759
336
0.809
0.748
0.826
0.759
0.791
0.741
720
0.896
0.794
0.919
0.803
0.886
0.790
S
96
0.446
0.482
0.476
0.508
0.460
0.487
192
0.462
0.491
0.511
0.528
0.505
0.511
336
0.521
0.525
0.576
0.568
0.569
0.545
720
0.717
0.625
0.795
0.685
0.764
0.641
AvgWins
85.0% 2.5% 12.5%
9 Exploratory Studies in Generalist Modeling
Generalist Codebooks. To further explore the capabilities of a generalist codebook data representation we train models that utilize a general codebook but dataset-specific transformer forecasters, e.g. a TOTEM VQVAE trained on multiple domains with a forecaster trained only on electricity, Table 8. We compare these mixed models to generalist and specialist models trained on the same domains. All models use the same the codebook hyperparameters (number of codewords $K=256$, compression factor $F=4$, code dimensionality $D=64$) as well as the forecaster transformer architecture to ensure a fair comparison.
Since we are evaluating the specialists, mixed-models, and generalist on in-domain test data one might expect that the TOTEM specialists will significantly outperform all models. Surprisingly this intuition is not correct. When comparing models trained using specialist codebooks to models trained using a single generalist codebook we find that generalist codebook models outperform specialist codebook models: 66.1% vs. 57.1%. Upon further inspection we find that the fully-generalist model (far right column Table 8) significantly outperforms the mixed-models (middle column Table 8) in traffic (T) and electricity (E). This dominant performance is puzzling until considering the training sizes.
The largest training set across domains belongs to traffic (T) at $10.2M$ training examples. In dataset T, the fully generalist models achieves 100% AvgWins . The second largest training set belongs to electricity (E) at $5.8M$ training examples, with 75% AvgWins for the fully-generalist model. Unfortunately there is a sharp drop off in training set sizes, with the rest of the data domains collectively comprising $1.6M$ training examples. These results evoke questions. For instance: does training on the smaller datasets act like form of regularization? Or: how does in-domain generalist performance scale with dataset size? We leave these exciting directions for future work. The generalist codebook’s performance across datasets highlights the potential of unified, discrete, token representations for in-domain evaluations.
Zero Shot Vignette: Training Size & Data Diversity. Here we further explore generalist and specialist zero-shot testing capabilites, Table 9. We take the two largest TOTEM specialist, traffic at $10.2M$ and electricity at $5.8M$ training examples, and test their zero-shot capabilities compared to the TOTEM generalist. We expect that the generalist will perform best as it was trained on the most data at $17.6M$ training examples as well as the most domains. We predict the generalist will be followed by TOTEM-traffic then TOTEM-electricity as they are both trained on only one domain but traffic has $4.4M$ more training examples than electricity. As expected the generalist outperforms both TOTEM-traffic and TOTEM-electricity with 85.0% AvgWins . However, curiously TOTEM-electricity outperforms TOTEM-traffic: 12.5% vs. 2.5% despite having $4.4M$ fewer training examples. Why is the smaller training set outperforming the larger training set? One possible explanation is that the electricity domain is more similar than the traffic domain to neuro, river, births, and sunspot. Another possible explanation comes from the raw time series dimensionality. Despite having fewer training examples, electricity has a higher number of raw time steps [^4] Raw time steps for all data. The train:val:test ratio is 7:1:2. compared to traffic: 26304 vs. 17544. However, traffic has a larger number of sensors: 862 vs. 321. This limited analysis suggests that a higher number of raw time steps is more valuable than more sensor readings. Untangling these possibilities and beginning to answer the questions: what is a unit of data in time series? And how does this unit scale as the time steps, sensors, and examples scale? are valuable future directions. The zero shot vignette has demonstrated the power of the token-enabled generalist over the traffic and electricity specialists, and has opened up exciting training size and data diversity questions.
10 Conclusions, Limitations & Future Work
We present TOTEM: a simple, performant tokenizer that works across domains and tasks thereby enabling generalist modeling. TOTEM demonstrates strong in-domain and zero-shot capabilities that match or outperform existing state-of-the-art approaches. Moving forward, an interesting limitation is that TOTEM does not support variable token lengths. Dynamic token lengths could potentially enhance unified data representations and further improve task performance. Other interesting directions include further investigating the relationship between generalist data representations, token length, data size, and domain diversity.
11 Impact Statement
There are no immediate ethical concerns that arise from our work. However, as with all data driven methods, certain societal consequences are important to be discussed, in this case surrounding time series modeling. A few are reported below:
Privacy Concerns. Time series data, especially when sourced from personal devices or applications, can contain sensitive information about individuals, e.g. for health domains. In this work, no time series were sourced from personal devices.
Misuse. Time series forecast models can be misused. For instance, if a model forecasts stock prices or market movements, it could be exploited for insider trading or other illegal financial activities. In this work, we are focused on domains pertinent to scientific disciplines.
Economic Impacts. Automated forecasts and decisions based on time series models can significantly impact industries and labor markets both positively and negatively. For instance, if a model can accurately predict weather patterns, it might affect farmers and their crop decisions, or if it can forecast energy consumption, it could impact the energy sector.
Appendix A Appendix
A.1 Codebook Visualizations.
Figure 3: TOTEM Codebooks. We visualize all 256 codes for the generalist (All), and three specialists (Traffic, Electricity, and Weather). The top row visualizes codes in the latent space, the bottom row visualizes codes in the decoded time space. We additionally highlight codeword pairs matched via low MSE between All-Traffic, All-Electricity, and All-Weather in the bottom row.
A.2 TOTEM Examples.
Figure 4: TOTEM Examples. In the top row we visualize four weather forecasts for Tin=96 and Tout=96. In the bottom row we visualize four ETTm2 imputations. In all cases the model input is in grey, the predictions are in blue, and the ground truth is in green.
A.3 Architecture Details.
VQVAE. For imputation, anomaly detection, and forecasting the VQVAE’s number of residual layers = 2, residual hidden size = 64, and block hidden size = 128 for all datasets. Each residual block has 2 non-causal, non-dilated 1D convolutional layers. The residual blocks are paired with additional non-causal, non-dilated 1D convolutional layers, where the number of additional layers is determined by the desired compression factor. See Table 10 for more hyperparameter details.
Table 10: VQVAE Hyperparameters (A) Imputation generalist (All) and specialists. (B) Anomaly detection generalist (All) and specialists. The anomaly %s for all of the zero shot datasets are 2%. (C) Forecasting generalist (All) and specialists.
A. Imputation.
Dataset LR Iter. BS # CW CW Dim. CF
All 1e-3 120000 8192 512 64 4
Elec. 1e-3 15000 8192 512 64 4
Weather 1e-3 15000 8192 512 64 4
ETTm1 1e-3 15000 8192 512 64 4
ETTm2 1e-3 15000 8192 512 64 4
ETTh1 1e-3 15000 8192 512 64 4
ETTh2 1e-3 15000 8192 512 64 4
B. Anomaly Detection.
Dataset LR Iter. BS # CW CW Dim. CF Anomaly %
All 1e-3 120000 4096 1024 64 4 Varies by test set.
SMD 1e-3 60000 4096 1024 64 4 0.5
MSL 1e-3 15000 4096 1024 64 4 2
PSM 1e-3 60000 4096 1024 64 4 1
SMAP 1e-3 15000 4096 1024 64 4 1
SWAT 1e-3 15000 4096 1024 64 4 1
C. Forecasting.
Dataset LR Iter. BS # CW CW Dim. CF
All 1e-3 15000 4096 256 64 4
Elec. 1e-3 15000 4096 256 64 4
Weather 1e-3 15000 4096 256 64 4
Traffic 1e-3 15000 4096 256 64 4
ETTm1 1e-3 15000 4096 256 64 4
ETTm2 1e-3 15000 4096 256 64 4
ETTh1 1e-3 15000 4096 256 64 4
ETTh2 1e-3 15000 4096 256 64 4
Downstream Forecaster. The downstream forecaster has two components the transformer encoder that intakes codes and outputs a normalized time forecast, and the feedforward neural network that takes in time and outputs predictions for the forecast’s mean and standard deviation. The downstream forecaster is a transformer encoder with a model dimension = 64, hidden dimension = 256, number of heads = 4, number of layers = 4. The transformer encoder applies a sin / cos positional embedding along the time dimension and applies its attention mechanism to each sensor independently. There is a single linear layer applied after the transformer encoder output. The feedforward neural network takes in the input time steps, and predicts the future’s mean and standard deviation.
A.4 Training Details.
In imputation, anomaly detection, and forecasting the VQVAE is trained with a learning rate of $0.001$ using the Adam optimizer, embedding dimension of 64, commitment cost of 0.25, and compression factor of 4; see Table 10 for more hyperparameters. In all tasks there is a global normalization, and local normalization (Kim et al., 2021); both are standard throughout prior work. In imputation we only leverage global normalization, in anomaly detection and forecasting we utilize both global and local normalization. In anomaly detection we evaluate the models we run, TOTEM and GPT2, with both local normalized data and non-local normalized data for each method and report whichever schema leads to the best performance. In forecasting the downstream model is a transformer encoder with 4 layers and 4 attention heads and a feed-forward hidden dimension of 256. We train using Adam with a base learning rate of $0.0001$ and a one cycle learning rate scheduler in accordance with (Nie et al., 2022) on A100s.
A.5 Imputation Results - Means and Std. Devs.
Table 11: TOTEM - Specialist Imputation ()normal-↓($\downarrow$)( ↓ )
Metric
MSE
MAE
W
12.5%
$0.028\pm 0.0000$
$0.046\pm 0.0006$
37.5%
$0.029\pm 0.0000$
$0.047\pm 0.0010$
50%
$0.031\pm 0.0006$
$0.048\pm 0.0015$
25%
$0.033\pm 0.0000$
$0.052\pm 0.0006$
E
12.5%
$0.054\pm 0.0006$
$0.154\pm 0.0015$
25%
$0.059\pm 0.0006$
$0.160\pm 0.0010$
37.5%
$0.067\pm 0.0006$
$0.169\pm 0.0012$
50%
$0.079\pm 0.0012$
$0.183\pm 0.0012$
m1
12.5%
$0.049\pm 0.0000$
$0.125\pm 0.0006$
25%
$0.052\pm 0.0006$
$0.128\pm 0.0006$
37.5%
$0.055\pm 0.0000$
$0.132\pm 0.0006$
50%
$0.061\pm 0.0006$
$0.139\pm 0.0006$
m2
12.5%
$0.016\pm 0.0006$
$0.078\pm 0.0010$
25%
$0.017\pm 0.0006$
$0.081\pm 0.0006$
37.5%
$0.018\pm 0.0000$
$0.084\pm 0.0006$
50%
$0.020\pm 0.0000$
$0.088\pm 0.0000$
h1
12.5%
$0.119\pm 0.0010$
$0.212\pm 0.0006$
25%
$0.127\pm 0.0015$
$0.220\pm 0.0006$
37.5%
$0.138\pm 0.0012$
$0.230\pm 0.0006$
50%
$0.157\pm 0.0006$
$0.247\pm 0.0010$
h2
12.5%
$0.040\pm 0.0006$
$0.129\pm 0.0017$
25%
$0.041\pm 0.0010$
$0.131\pm 0.0012$
37.5%
$0.043\pm 0.0006$
$0.136\pm 0.0006$
50%
$0.047\pm 0.0006$
$0.142\pm 0.0012$
Table 12: TOTEM - Generalist Imputation ()normal-↓($\downarrow$)( ↓ )
Metric
MSE
MAE
W
12.5%
$0.029\pm 0.0012$
$0.060\pm 0.0047$
25%
$0.030\pm 0.0006$
$0.060\pm 0.0047$
37.5%
$0.032\pm 0.0006$
$0.062\pm 0.0030$
50%
$0.036\pm 0.0006$
$0.067\pm 0.0036$
E
12.5%
$0.065\pm 0.0020$
$0.171\pm 0.0032$
25%
$0.071\pm 0.0015$
$0.179\pm 0.0031$
37.5%
$0.080\pm 0.0025$
$0.189\pm 0.0032$
50%
$0.095\pm 0.0026$
$0.205\pm 0.0032$
m1
12.5%
$0.041\pm 0.0006$
$0.132\pm 0.0015$
25%
$0.044\pm 0.0000$
$0.135\pm 0.0010$
37.5%
$0.048\pm 0.0006$
$0.139\pm 0.0040$
50%
$0.058\pm 0.0010$
$0.152\pm 0.0000$
m2
12.5%
$0.040\pm 0.0020$
$0.125\pm 0.0067$
25%
$0.041\pm 0.0015$
$0.126\pm 0.0058$
37.5%
$0.043\pm 0.0015$
$0.129\pm 0.0049$
50%
$0.048\pm 0.0010$
$0.136\pm 0.0038$
h1
12.5%
$0.100\pm 0.0049$
$0.201\pm 0.0049$
25%
$0.108\pm 0.0049$
$0.209\pm 0.0038$
37.5%
$0.122\pm 0.0064$
$0.220\pm 0.0044$
50%
$0.144\pm 0.0078$
$0.237\pm 0.0049$
h2
12.5%
$0.075\pm 0.0012$
$0.175\pm 0.0053$
25%
$0.076\pm 0.0006$
$0.177\pm 0.0036$
37.5%
$0.093\pm 0.0222$
$0.195\pm 0.0200$
50%
$0.089\pm 0.0010$
$0.192\pm 0.0035$
Zero-Shot
N2
12.5%
$0.029\pm 0.0015$
$0.120\pm 0.0045$
25%
$0.033\pm 0.0010$
$0.127\pm 0.0035$
37.5%
$0.041\pm 0.0006$
$0.139\pm 0.0025$
50%
$0.056\pm 0.0006$
$0.160\pm 0.0012$
N5
12.5%
$0.017\pm 0.0010$
$0.085\pm 0.0030$
25%
$0.019\pm 0.0010$
$0.090\pm 0.0030$
37.5%
$0.022\pm 0.0006$
$0.098\pm 0.0025$
50%
$0.029\pm 0.0006$
$0.110\pm 0.0025$
R
12.5%
$0.071\pm 0.0070$
$0.109\pm 0.0040$
25%
$0.087\pm 0.0064$
$0.117\pm 0.0031$
37.5%
$0.112\pm 0.0050$
$0.129\pm 0.0035$
50%
$0.148\pm 0.0032$
$0.147\pm 0.0023$
B
12.5%
$0.632\pm 0.0087$
$0.642\pm 0.0068$
25%
$0.693\pm 0.0070$
$0.665\pm 0.0047$
37.5%
$0.761\pm 0.0055$
$0.692\pm 0.0023$
50%
$0.827\pm 0.0044$
$0.718\pm 0.0000$
S
12.5%
$0.057\pm 0.0012$
$0.160\pm 0.0023$
25%
$0.061\pm 0.0006$
$0.168\pm 0.0021$
37.5%
$0.069\pm 0.0006$
$0.178\pm 0.0021$
50%
$0.082\pm 0.0010$
$0.193\pm 0.0015$
Table 13: GPT2 - Generalist Imputation ()normal-↓($\downarrow$)( ↓ )
Metric
MSE
MAE
W
12.5%
$0.029\pm 0.0000$
$0.045\pm 0.0006$
25%
$0.033\pm 0.0006$
$0.048\pm 0.0006$
37.5%
$0.037\pm 0.0006$
$0.054\pm 0.0012$
50%
$0.043\pm 0.0012$
$0.061\pm 0.0017$
E
12.5%
$0.008\pm 0.0020$
$0.186\pm 0.0035$
25%
$0.091\pm 0.0020$
$0.197\pm 0.0025$
37.5%
$0.108\pm 0.0021$
$0.213\pm 0.0026$
50%
$0.132\pm 0.0026$
$0.236\pm 0.0026$
m1
12.5%
$0.052\pm 0.0012$
$0.141\pm 0.0016$
25%
$0.065\pm 0.0021$
$0.154\pm 0.0021$
37.5%
$0.085\pm 0.0038$
$0.171\pm 0.0026$
50%
$0.117\pm 0.0052$
$0.196\pm 0.0026$
m2
12.5%
$0.029\pm 0.0000$
$0.095\pm 0.0006$
25%
$0.033\pm 0.0006$
$0.101\pm 0.0006$
37.5%
$0.038\pm 0.0006$
$0.110\pm 0.0012$
50%
$0.045\pm 0.0006$
$0.121\pm 0.0012$
h1
12.5%
$0.113\pm 0.0012$
$0.217\pm 0.0021$
25%
$0.131\pm 0.0010$
$0.231\pm 0.0015$
37.5%
$0.153\pm 0.0012$
$0.247\pm 0.0017$
50%
$0.182\pm 0.0006$
$0.266\pm 0.0012$
h2
12.5%
$0.067\pm 0.0010$
$0.155\pm 0.0015$
25%
$0.071\pm 0.0006$
$0.160\pm 0.0015$
37.5%
$0.077\pm 0.0010$
$0.167\pm 0.0015$
50%
$0.086\pm 0.0032$
$0.179\pm 0.0038$
Zero-Shot
N2
12.5%
$0.047\pm 0.0006$
$0.145\pm 0.0015$
25%
$0.064\pm 0.0017$
$0.164\pm 0.0015$
37.5%
$0.090\pm 0.0036$
$0.191\pm 0.0032$
50%
$0.131\pm 0.0051$
$0.228\pm 0.0044$
N5
12.5%
$0.021\pm 0.0006$
$0.095\pm 0.0012$
25%
$0.028\pm 0.0006$
$0.107\pm 0.0010$
37.5%
$0.039\pm 0.0015$
$0.123\pm 0.0015$
50%
$0.055\pm 0.0015$
$0.145\pm 0.0023$
R
12.5%
$0.093\pm 0.0010$
$0.119\pm 0.0015$
25%
$0.125\pm 0.0006$
$0.134\pm 0.0026$
37.5%
$0.167\pm 0.0021$
$0.154\pm 0.0042$
50%
$0.220\pm 0.0045$
$0.182\pm 0.0057$
B
12.5%
$0.392\pm 0.0064$
$0.496\pm 0.0023$
25%
$0.444\pm 0.0071$
$0.523\pm 0.0029$
37.5%
$0.498\pm 0.0080$
$0.553\pm 0.0023$
50%
$0.591\pm 0.0700$
$0.599\pm 0.0275$
s
12.5%
$0.070\pm 0.0012$
$0.173\pm 0.0017$
25%
$0.084\pm 0.0010$
$0.189\pm 0.0015$
37.5%
$0.103\pm 0.0010$
$0.209\pm 0.0021$
50%
$0.128\pm 0.0015$
$0.234\pm 0.0021$
A.6 Anomaly Detection Results - Means and Std Devs.
Table 14: TOTEM - Specialist Anomaly Detection ()normal-↑($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7962\pm 0.0137$
MSL
$0.8258\pm 0.0052$
SMAP
$0.9402\pm 0.0008$
SWAT
$0.9427\pm 0.0006$
PSM
$0.9587\pm 0.0008$
R
SMD
$0.7606\pm 0.0207$
MSL
$0.8285\pm 0.0071$
SMAP
$0.9404\pm 0.0013$
SWAT
$0.9591\pm 0.0012$
PSM
$0.9421\pm 0.0004$
P
SMD
$0.8354\pm 0.0054$
MSL
$0.8232\pm 0.0033$
SMAP
$0.9400\pm 0.0004$
SWAT
$0.9268\pm 0.0003$
PSM
$0.9758\pm 0.0012$
Table 15: TOTEM - Generalist Anomaly Detection ()normal-↑($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7864\pm 0.0386$
MSL
$0.8329\pm 0.0020$
SMAP
$0.9251\pm 0.0014$
SWAT
$0.9437\pm 0.0005$
PSM
$0.9578\pm 0.0002$
N2
$0.5129\pm 0.0397$
N5
$0.5128\pm 0.0390$
R
$0.4939\pm 0.0625$
B
$0.4915\pm 0.0229$
S
$0.5217\pm 0.0418$
R
SMD
$0.7207\pm 0.0565$
MSL
$0.8296\pm 0.0046$
SMAP
$0.9148\pm 0.0020$
SWAT
$0.9613\pm 0.0010$
PSM
$0.9390\pm 0.0004$
N2
$0.7688\pm 0.0594$
N5
$0.7684\pm 0.0582$
R
$0.7049\pm 0.0825$
B
$0.7371\pm 0.0340$
S
$0.7736\pm 0.0581$
P
SMD
$0.8666\pm 0.0114$
MSL
$0.8364\pm 0.0014$
SMAP
$0.9356\pm 0.0009$
SWAT
$0.9268\pm 0.0001$
PSM
$0.9774\pm 0.0002$
N2
$0.3849\pm 0.0299$
N5
$0.3848\pm 0.0294$
R
$0.3802\pm 0.0502$
B
$0.3686\pm 0.0172$
S
$0.3935\pm 0.0325$
Table 16: GPT2 - Generalist Anomaly Detection ()normal-↑($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7973\pm 0.0326$
MSL
$0.8017\pm 0.0205$
SMAP
$0.6705\pm 0.0041$
SWAT
$0.8962\pm 0.0016$
PSM
$0.9047\pm 0.0759$
N2
$0.3902\pm 0.0596$
N5
$0.4219\pm 0.0047$
R
$0.3614\pm 0.0204$
B
$0.2081\pm 0.0462$
S
$0.3812\pm 0.0621$
R
SMD
$0.7342\pm 0.0559$
MSL
$0.7848\pm 0.0277$
SMAP
$0.5342\pm 0.0051$
SWAT
$0.8753\pm 0.0033$
PSM
$0.8776\pm 0.0624$
N2
$0.3369\pm 0.0592$
N5
$0.3677\pm 0.0498$
R
$0.2966\pm 0.0218$
B
$0.1767\pm 0.0426$
S
$0.3183\pm 0.0648$
P
SMD
$0.8744\pm 0.0029$
MSL
$0.8195\pm 0.0130$
SMAP
$0.9001\pm 0.0007$
SWAT
$0.9183\pm 0.0006$
PSM
$0.9339\pm 0.0925$
N2
$0.4643\pm 0.0561$
N5
$0.4958\pm 0.0396$
R
$0.4630\pm 0.0139$
B
$0.2533\pm 0.0498$
S
$0.4772\pm 0.5000$
A.7 Forecasting Results - Means and Std Devs.
Table 17: TOTEM - Specialist Forecasting ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.165\pm 0.0015$
$0.208\pm 0.0012$
192
$0.207\pm 0.0006$
$0.250\pm 0.0012$
336
$0.257\pm 0.0002$
$0.291\pm 0.0006$
720
$0.326\pm 0.0035$
$0.340\pm 0.0023$
E
96
$0.178\pm 0.0015$
$0.263\pm 0.0010$
192
$0.187\pm 0.0015$
$0.272\pm 0.0015$
336
$0.199\pm 0.0012$
$0.285\pm 0.0012$
720
$0.236\pm 0.0035$
$0.318\pm 0.0031$
T
96
$0.523\pm 0.0010$
$0.303\pm 0.0006$
192
$0.530\pm 0.0030$
$0.303\pm 0.0017$
336
$0.549\pm 0.0017$
$0.311\pm 0.0021$
720
$0.598\pm 0.0095$
$0.331\pm 0.0062$
m1
96
$0.320\pm 0.0006$
$0.347\pm 0.0006$
192
$0.379\pm 0.0017$
$0.382\pm 0.0012$
336
$0.406\pm 0.0040$
$0.402\pm 0.0026$
720
$0.471\pm 0.0006$
$0.438\pm 0.0010$
m2
96
$0.176\pm 0.0006$
$0.253\pm 0.0010$
192
$0.247\pm 0.0012$
$0.302\pm 0.0015$
336
$0.317\pm 0.0046$
$0.348\pm 0.0031$
720
$0.426\pm 0.0085$
$0.410\pm 0.0062$
h1
96
$0.380\pm 0.0006$
$0.394\pm 0.0000$
192
$0.434\pm 0.0010$
$0.427\pm 0.0006$
336
$0.490\pm 0.0023$
$0.459\pm 0.0015$
720
$0.539\pm 0.0031$
$0.513\pm 0.0020$
h2
96
$0.293\pm 0.0015$
$0.338\pm 0.0006$
192
$0.375\pm 0.0031$
$0.390\pm 0.0026$
336
$0.422\pm 0.0046$
$0.431\pm 0.0031$
720
$0.610\pm 0.0095$
$0.567\pm 0.0081$
Table 18: GPT2 - Specialist Forecasting, Lookback of 96 ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.184\pm 0.0013$
$0.224\pm 0.0014$
192
$0.231\pm 0.0012$
$0.263\pm 0.0009$
336
$0.285\pm 0.0015$
$0.302\pm 0.0013$
720
$0.362\pm 0.0016$
$0.351\pm 0.0008$
E
96
$0.186\pm 0.0004$
$0.272\pm 0.0005$
192
$0.190\pm 0.0007$
$0.278\pm 0.0008$
336
$0.204\pm 0.0003$
$0.291\pm 0.0005$
720
$0.245\pm 0.0012$
$0.324\pm 0.0014$
T
96
$0.471\pm 0.0016$
$0.311\pm 0.0016$
192
$0.479\pm 0.0017$
$0.312\pm 0.0010$
336
$0.490\pm 0.0009$
$0.317\pm 0.0010$
720
$0.524\pm 0.0019$
$0.336\pm 0.0018$
m1
96
$0.328\pm 0.0022$
$0.363\pm 0.0014$
192
$0.368\pm 0.0006$
$0.382\pm 0.0004$
336
$0.400\pm 0.0013$
$0.404\pm 0.0011$
720
$0.462\pm 0.0010$
$0.440\pm 0.0009$
m2
96
$0.178\pm 0.0000$
$0.263\pm 0.0000$
192
$0.245\pm 0.0000$
$0.307\pm 0.0000$
336
$0.307\pm 0.0000$
$0.346\pm 0.0000$
720
$0.410\pm 0.0000$
$0.409\pm 0.0000$
h1
96
$0.379\pm 0.0032$
$0.397\pm 0.0007$
192
$0.438\pm 0.0037$
$0.427\pm 0.0004$
336
$0.474\pm 0.0045$
$0.448\pm 0.0004$
720
$0.496\pm 0.0066$
$0.475\pm 0.0033$
h2
96
$0.295\pm 0.0000$
$0.348\pm 0.0000$
192
$0.384\pm 0.0000$
$0.402\pm 0.0000$
336
$0.418\pm 0.0000$
$0.432\pm 0.0000$
720
$0.423\pm 0.0000$
$0.446\pm 0.0000$
Table 19: TOTEM - Generalist and Zero-Shot Forecasting ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.172\pm 0.0010$
$0.216\pm 0.0006$
192
$0.217\pm 0.0006$
$0.256\pm 0.0006$
336
$0.266\pm 0.0015$
$0.295\pm 0.0015$
720
$0.334\pm 0.0010$
$0.342\pm 0.0012$
E
96
$0.179\pm 0.0006$
$0.264\pm 0.0012$
192
$0.181\pm 0.0006$
$0.267\pm 0.0000$
336
$0.196\pm 0.0020$
$0.283\pm 0.0015$
720
$0.230\pm 0.0035$
$0.314\pm 0.0029$
T
96
$0.507\pm 0.0020$
$0.284\pm 0.0006$
192
$0.511\pm 0.0030$
$0.282\pm 0.0006$
336
$0.535\pm 0.0076$
$0.292\pm 0.0012$
720
$0.580\pm 0.0046$
$0.309\pm 0.0006$
m1
96
$0.374\pm 0.0000$
$0.384\pm 0.0006$
192
$0.400\pm 0.0015$
$0.399\pm 0.0023$
336
$0.432\pm 0.0040$
$0.424\pm 0.0015$
720
$0.487\pm 0.0081$
$0.460\pm 0.0017$
m2
96
$0.198\pm 0.0006$
$0.275\pm 0.0012$
192
$0.266\pm 0.0035$
$0.319\pm 0.0021$
336
$0.365\pm 0.0115$
$0.377\pm 0.0038$
720
$0.588\pm 0.0699$
$0.511\pm 0.0281$
h1
96
$0.382\pm 0.0364$
$0.404\pm 0.0012$
192
$0.463\pm 0.0025$
$0.435\pm 0.0006$
336
$0.507\pm 0.0025$
$0.463\pm 0.0010$
720
$0.517\pm 0.0010$
$0.500\pm 0.0017$
h2
96
$0.307\pm 0.0012$
$0.345\pm 0.0015$
192
$0.406\pm 0.0038$
$0.403\pm 0.0023$
336
$0.505\pm 0.0114$
$0.460\pm 0.0035$
720
$0.661\pm 0.0514$
$0.557\pm 0.0215$
Zero-Shot
N2
96
$1.138\pm 0.0032$
$0.777\pm 0.0012$
192
$1.149\pm 0.0026$
$0.785\pm 0.0012$
336
$1.092\pm 0.0062$
$0.770\pm 0.0026$
720
$1.045\pm 0.0040$
$0.754\pm 0.0023$
N5
96
$0.483\pm 0.0012$
$0.484\pm 0.0012$
192
$0.495\pm 0.0021$
$0.491\pm 0.0015$
336
$0.468\pm 0.0035$
$0.483\pm 0.0029$
720
$0.451\pm 0.0023$
$0.477\pm 0.0023$
R
96
$1.120\pm 0.0081$
$0.582\pm 0.0036$
192
$1.242\pm 0.0151$
$0.635\pm 0.0074$
336
$1.237\pm 0.0153$
$0.626\pm 0.0076$
720
$1.182\pm 0.0151$
$0.604\pm 0.0050$
B
96
$0.805\pm 0.0070$
$0.739\pm 0.0035$
192
$0.836\pm 0.0040$
$0.752\pm 0.0021$
336
$0.809\pm 0.0038$
$0.748\pm 0.0021$
720
$0.896\pm 0.0137$
$0.794\pm 0.0085$
S
96
$0.446\pm 0.0032$
$0.482\pm 0.0017$
192
$0.462\pm 0.0015$
$0.491\pm 0.0010$
336
$0.521\pm 0.0122$
$0.525\pm 0.0068$
720
$0.717\pm 0.0096$
$0.625\pm 0.0040$
Table 20: GPT2 - Generalist and Zero-Shot Forecasting ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.201\pm 0.0017$
$0.237\pm 0.0012$
192
$0.247\pm 0.0020$
$0.275\pm 0.0015$
336
$0.298\pm 0.0006$
$0.311\pm 0.0006$
720
$0.372\pm 0.0010$
$0.360\pm 0.0006$
E
96
$0.194\pm 0.0012$
$0.278\pm 0.0021$
192
$0.199\pm 0.0006$
$0.284\pm 0.0006$
336
$0.214\pm 0.0012$
$0.300\pm 0.0015$
720
$0.255\pm 0.0006$
$0.331\pm 0.0012$
T
96
$0.484\pm 0.0046$
$0.320\pm 0.0042$
192
$0.488\pm 0.0006$
$0.320\pm 0.0006$
336
$0.502\pm 0.0020$
$0.326\pm 0.0021$
720
$0.534\pm 0.0021$
$0.343\pm 0.0021$
m1
96
$0.487\pm 0.0106$
$0.468\pm 0.0035$
192
$0.516\pm 0.0071$
$0.480\pm 0.0021$
336
$0.548\pm 0.0015$
$0.499\pm 0.0015$
720
$0.581\pm 0.0031$
$0.511\pm 0.0012$
m2
96
$0.243\pm 0.0021$
$0.315\pm 0.0021$
192
$0.297\pm 0.0012$
$0.346\pm 0.0010$
336
$0.349\pm 0.0025$
$0.376\pm 0.0020$
720
$0.439\pm 0.0010$
$0.423\pm 0.0010$
h1
96
$0.421\pm 0.0058$
$0.408\pm 0.0010$
192
$0.480\pm 0.0026$
$0.436\pm 0.0020$
336
$0.518\pm 0.0161$
$0.453\pm 0.0070$
720
$0.517\pm 0.0036$
$0.467\pm 0.0035$
h2
96
$0.298\pm 0.0090$
$0.343\pm 0.0049$
192
$0.381\pm 0.0153$
$0.392\pm 0.0072$
336
$0.406\pm 0.0271$
$0.419\pm 0.0144$
720
$0.423\pm 0.0078$
$0.438\pm 0.0051$
Zero-Shot
N2
96
$1.332\pm 0.0012$
$0.830\pm 0.0010$
192
$1.416\pm 0.0080$
$0.863\pm 0.0025$
336
$1.358\pm 0.0123$
$0.851\pm 0.0042$
720
$1.308\pm 0.0026$
$0.840\pm 0.0010$
N5
96
$0.528\pm 0.0006$
$0.499\pm 0.0010$
192
$0.578\pm 0.0015$
$0.524\pm 0.0006$
336
$0.548\pm 0.0040$
$0.515\pm 0.0015$
720
$0.537\pm 0.0006$
$0.511\pm 0.0006$
R
96
$1.465\pm 0.0185$
$0.725\pm 0.0031$
192
$1.638\pm 0.0280$
$0.785\pm 0.0078$
336
$1.601\pm 0.0244$
$0.769\pm 0.0060$
720
$1.552\pm 0.0110$
$0.760\pm 0.0035$
B
96
$0.838\pm 0.0149$
$0.762\pm 0.0071$
192
$0.837\pm 0.0095$
$0.752\pm 0.0040$
336
$0.792\pm 0.0104$
$0.738\pm 0.0050$
720
$0.927\pm 0.0066$
$0.806\pm 0.0038$
S
96
$0.443\pm 0.0010$
$0.478\pm 0.0006$
192
$0.481\pm 0.0006$
$0.499\pm 0.0006$
336
$0.541\pm 0.0010$
$0.533\pm 0.0006$
720
$0.773\pm 0.0020$
$0.643\pm 0.0010$
A.8 Additional Ablations
Table 21: TimeTOTEM Ablation - Specialist Forecasting
Mean $\pm$ Std
MSE
MAE
W
96
$0.164\pm 0.0006$
$0.209\pm 0.0006$
192
$0.209\pm 0.0017$
$0.251\pm 0.0023$
336
$0.261\pm 0.0012$
$0.293\pm 0.0017$
720
$0.332\pm 0.0023$
$0.340\pm 0.0006$
E
96
$0.179\pm 0.0015$
$0.262\pm 0.0015$
192
$0.185\pm 0.0006$
$0.269\pm 0.0000$
336
$0.204\pm 0.0055$
$0.289\pm 0.0061$
720
$0.244\pm 0.0040$
$0.325\pm 0.0036$
T
96
$0.528\pm 0.0081$
$0.310\pm 0.0092$
192
$0.500\pm 0.0606$
$0.349\pm 0.0699$
336
$0.531\pm 0.0424$
$0.365\pm 0.0852$
720
$0.578\pm 0.0361$
$0.398\pm 0.1103$
m1
96
$0.326\pm 0.0006$
$0.355\pm 0.0006$
192
$0.377\pm 0.0023$
$0.386\pm 0.0012$
336
$0.409\pm 0.0006$
$0.409\pm 0.0006$
720
$0.469\pm 0.0015$
$0.441\pm 0.0000$
m2
96
$0.176\pm 0.0010$
$0.254\pm 0.0006$
192
$0.247\pm 0.0031$
$0.303\pm 0.0026$
336
$0.318\pm 0.0006$
$0.350\pm 0.0021$
720
$0.419\pm 0.0067$
$0.411\pm 0.0044$
h1
96
$0.377\pm 0.0010$
$0.395\pm 0.0006$
192
$0.428\pm 0.0015$
$0.428\pm 0.0015$
336
$0.480\pm 0.0021$
$0.462\pm 0.0012$
720
$0.530\pm 0.0110$
$0.522\pm 0.0108$
h2
96
$0.294\pm 0.0021$
$0.338\pm 0.0010$
192
$0.373\pm 0.0023$
$0.389\pm 0.0032$
336
$0.423\pm 0.0031$
$0.433\pm 0.0025$
720
$0.591\pm 0.0145$
$0.556\pm 0.0051$
Table 22: TimeTOTEM Ablation - Generalist and Zero-Shot Forecasting
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.173\pm 0.0012$
$0.218\pm 0.0006$
192
$0.218\pm 0.0006$
$0.261\pm 0.0006$
336
$0.267\pm 0.0006$
$0.299\pm 0.0006$
720
$0.337\pm 0.0010$
$0.347\pm 0.0006$
E
96
$0.183\pm 0.0012$
$0.267\pm 0.0012$
192
$0.189\pm 0.0006$
$0.275\pm 0.0000$
336
$0.204\pm 0.0010$
$0.291\pm 0.0010$
720
$0.242\pm 0.0006$
$0.325\pm 0.0006$
T
96
$0.517\pm 0.0000$
$0.293\pm 0.0029$
192
$0.526\pm 0.0030$
$0.296\pm 0.0006$
336
$0.552\pm 0.0015$
$0.304\pm 0.0015$
720
$0.602\pm 0.0046$
$0.326\pm 0.0015$
m1
96
$0.428\pm 0.0090$
$0.420\pm 0.0040$
192
$0.438\pm 0.0015$
$0.427\pm 0.0010$
336
$0.469\pm 0.0062$
$0.447\pm 0.0042$
720
$0.546\pm 0.0081$
$0.493\pm 0.0017$
m2
96
$0.207\pm 0.0015$
$0.286\pm 0.0020$
192
$0.269\pm 0.0015$
$0.325\pm 0.0010$
336
$0.358\pm 0.0199$
$0.377\pm 0.0091$
720
$0.521\pm 0.0165$
$0.482\pm 0.0026$
h1
96
$0.401\pm 0.0006$
$0.410\pm 0.0006$
192
$0.453\pm 0.0010$
$0.441\pm 0.0010$
336
$0.496\pm 0.0017$
$0.468\pm 0.0006$
720
$0.518\pm 0.0020$
$0.510\pm 0.0017$
h2
96
$0.305\pm 0.0006$
$0.346\pm 0.0006$
192
$0.396\pm 0.0015$
$0.402\pm 0.0001$
336
$0.492\pm 0.0310$
$0.458\pm 0.0131$
720
$0.599\pm 0.0105$
$0.531\pm 0.0026$
N2
96
$1.127\pm 0.0017$
$0.773\pm 0.0006$
192
$1.169\pm 0.0032$
$0.793\pm 0.0010$
336
$1.115\pm 0.0010$
$0.780\pm 0.0006$
720
$1.070\pm 0.0035$
$0.766\pm 0.0010$
N5
96
$0.481\pm 0.0015$
$0.483\pm 0.0006$
192
$0.508\pm 0.0012$
$0.500\pm 0.0000$
336
$0.481\pm 0.0006$
$0.491\pm 0.0006$
720
$0.467\pm 0.0010$
$0.488\pm 0.0010$
R
96
$1.102\pm 0.0031$
$0.578\pm 0.0021$
192
$1.207\pm 0.0036$
$0.628\pm 0.0017$
336
$1.190\pm 0.0021$
$0.613\pm 0.0010$
720
$1.149\pm 0.0017$
$0.596\pm 0.0020$
B
96
$0.825\pm 0.0079$
$0.751\pm 0.0076$
192
$0.847\pm 0.0021$
$0.761\pm 0.0012$
336
$0.831\pm 0.0066$
$0.764\pm 0.0042$
720
$0.928\pm 0.0131$
$0.813\pm 0.0050$
S
96
$0.446\pm 0.0015$
$0.481\pm 0.0010$
192
$0.478\pm 0.0015$
$0.499\pm 0.0000$
336
$0.535\pm 0.0012$
$0.532\pm 0.0006$
720
$0.736\pm 0.0025$
$0.631\pm 0.0006$
Table 23: Detailed Codebook Ablation ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
$K$
MSE
MAE
All
$32$
$0.0451\pm 0.0014$
$0.1460\pm 0.0030$
$256$
$0.0192\pm 0.0003$
$0.0937\pm 0.0007$
$512$
$0.0184\pm 0.0025$
$0.0913\pm 0.0062$
W
$32$
$0.0393\pm 0.0005$
$0.1122\pm 0.0064$
$256$
$0.0161\pm 0.0004$
$0.0673\pm 0.0011$
$512$
$0.0128\pm 0.0011$
$0.0607\pm 0.0032$
E
$32$
$0.0463\pm 0.0007$
$0.1520\pm 0.0016$
$256$
$0.0209\pm 0.0012$
$0.1027\pm 0.0029$
$512$
$0.0152\pm 0.0005$
$0.0878\pm 0.0014$
T
$32$
$0.0312\pm 0.0007$
$0.1204\pm 0.0008$
$256$
$0.0120\pm 0.0003$
$0.0749\pm 0.0007$
$512$
$0.0101\pm 0.0012$
$0.0685\pm 0.0044$
A.9 Exploratory Results
Table 24: Mixed Models - Forecasting ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
W
96
$0.164\pm 0.0010$
$0.208\pm 0.0012$
192
$0.208\pm 0.0010$
$0.251\pm 0.0015$
336
$0.258\pm 0.0012$
$0.290\pm 0.0015$
720
$0.329\pm 0.0021$
$0.338\pm 0.0015$
E
96
$0.178\pm 0.0006$
$0.263\pm 0.0010$
192
$0.187\pm 0.0021$
$0.273\pm 0.0017$
336
$0.199\pm 0.0012$
$0.285\pm 0.0017$
720
$0.238\pm 0.0012$
$0.320\pm 0.0012$
T
96
$0.521\pm 0.0010$
$0.301\pm 0.0010$
192
$0.530\pm 0.0023$
$0.303\pm 0.0012$
336
$0.555\pm 0.0080$
$0.313\pm 0.0072$
720
$0.605\pm 0.0097$
$0.337\pm 0.0075$
m1
96
$0.328\pm 0.0036$
$0.352\pm 0.0006$
192
$0.377\pm 0.0021$
$0.383\pm 0.0012$
336
$0.408\pm 0.0035$
$0.404\pm 0.0021$
720
$0.470\pm 0.0035$
$0.440\pm 0.0021$
m2
96
$0.175\pm 0.0006$
$0.253\pm 0.0010$
192
$0.247\pm 0.0006$
$0.302\pm 0.0010$
336
$0.318\pm 0.0006$
$0.348\pm 0.0031$
720
$0.427\pm 0.0012$
$0.410\pm 0.0067$
h1
96
$0.382\pm 0.0025$
$0.395\pm 0.0015$
192
$0.437\pm 0.0012$
$0.427\pm 0.0006$
336
$0.490\pm 0.0015$
$0.460\pm 0.0021$
720
$0.536\pm 0.0031$
$0.512\pm 0.0032$
h2
96
$0.294\pm 0.0010$
$0.339\pm 0.0012$
192
$0.375\pm 0.0025$
$0.391\pm 0.0023$
336
$0.421\pm 0.0050$
$0.431\pm 0.0031$
720
$0.610\pm 0.0089$
$0.567\pm 0.0075$
Table 25: Traffic Only - Specialist Zero-Shot Performance ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
N2
96
$1.194\pm 0.0062$
$0.798\pm 0.0020$
192
$1.218\pm 0.0074$
$0.808\pm 0.0023$
336
$1.190\pm 0.0153$
$0.804\pm 0.0052$
720
$1.117\pm 0.0137$
$0.784\pm 0.0056$
N5
96
$0.515\pm 0.0026$
$0.505\pm 0.0012$
192
$0.535\pm 0.0051$
$0.514\pm 0.0028$
336
$0.524\pm 0.0071$
$0.513\pm 0.0030$
720
$0.500\pm 0.0064$
$0.507\pm 0.0032$
R
96
$1.171\pm 0.0023$
$0.635\pm 0.0019$
192
$1.273\pm 0.0090$
$0.673\pm 0.0042$
336
$1.232\pm 0.0055$
$0.653\pm 0.0022$
720
$1.198\pm 0.0057$
$0.642\pm 0.0041$
B
96
$0.812\pm 0.0037$
$0.749\pm 0.0025$
192
$0.858\pm 0.0025$
$0.767\pm 0.0015$
336
$0.826\pm 0.0041$
$0.759\pm 0.0030$
720
$0.919\pm 0.0063$
$0.803\pm 0.0037$
S
96
$0.476\pm 0.0012$
$0.508\pm 0.0012$
192
$0.511\pm 0.0005$
$0.528\pm 0.0005$
336
$0.576\pm 0.0024$
$0.568\pm 0.0009$
720
$0.795\pm 0.0017$
$0.685\pm 0.0012$
Table 26: Electricity Only - Specialist Zero-Shot Performance ()normal-↓($\downarrow$)( ↓ )
Mean $\pm$ Std
Metric
MSE
MAE
N2
96
$1.193\pm 0.0059$
$0.802\pm 0.0020$
192
$1.300\pm 0.0016$
$0.845\pm 0.0003$
336
$1.260\pm 0.0162$
$0.837\pm 0.0055$
720
$1.234\pm 0.0054$
$0.832\pm 0.0016$
N5
96
$0.489\pm 0.0024$
$0.490\pm 0.0011$
192
$0.555\pm 0.0012$
$0.527\pm 0.0007$
336
$0.538\pm 0.0064$
$0.525\pm 0.0033$
720
$0.533\pm 0.0010$
$0.527\pm 0.0006$
R
96
$1.141\pm 0.0056$
$0.579\pm 0.0028$
192
$1.297\pm 0.0162$
$0.652\pm 0.0079$
336
$1.247\pm 0.0108$
$0.628\pm 0.0059$
720
$1.236\pm 0.0053$
$0.633\pm 0.0070$
B
96
$0.820\pm 0.0065$
$0.756\pm 0.0034$
192
$0.843\pm 0.0042$
$0.759\pm 0.0022$
336
$0.791\pm 0.0023$
$0.741\pm 0.0019$
720
$0.886\pm 0.0059$
$0.790\pm 0.0020$
S
96
$0.460\pm 0.0017$
$0.487\pm 0.0010$
192
$0.505\pm 0.0017$
$0.511\pm 0.0008$
336
$0.569\pm 0.0020$
$0.545\pm 0.0011$
720
$0.764\pm 0.0046$
$0.641\pm 0.0014$